Can you please show me how to solve the system by addition?
x+y=9
x-y=9
Thank you

Found 2 solutions by Nate, stanbon:
Answer by Nate(3500)

x + x + y - y = 9 + 9
2x = 18
x = 9
plug:
9 - y = 9
y = 0

Answer by stanbon(75887)

1st: x+y=9
2nd: x-y=9
Add the two equations to get:
3rd: 2x=18
x=9
Substutute that into 1st to solve for y, as follows:
9+y=9
y=0
Check the answer in 2nd.
x-y=9
9-0=9
9=9
Cheers,
Stan H.